# Unbounded entropy production and violent fragmentation in long-range interacting super-Tonks-Girardeau systems

### v1

@@ VII Conclusions and Outlook
The present work addresses the nonequilibrium quench dynamics in a long-range interacting system, albeit with a different quench protocol. We investigated a generalized long-range interacting Bose gas where the strong repulsive interactions are suddenly quenched to strong attractive values, following the seminal observation of the exotic sTG gas in Haller’s experiment [62]. We presented a numerical study of the many-body dynamics for a wide range of interactions for both bosonic and fermionic systems. The usage of ab initio many-body methods facilitates a complete understanding of the exotic sTG phase in terms of fragmentation, delocalization and entropy production. We observed that the dynamics exhibited by the TG to sTG quench reveals new insights into the structure of quantum state relaxation.
Although the system eventually relaxes to a maximal entropy state, it passes through very complex phases of violent fragmentation and chaotic delocalization. The entropy production is determined by the number of configurations in the Hilbert space, which are saturated for all the numbers of orbitals that can be employed with current computational capabilities. The observation of unbounded entropy and chaotic spreading of the sTG gas in the available Hilbert space strongly suggests a behavior akin to that of a classical gas. We presented different measures of entropy for our discussion. However, we chose the occupation entropy as an order parameter to produce a parameter-time diagram that illustrates the relaxation process as a function of interaction length, covering the entire range of $4.00.5<\alpha<$ 4.0. We found an exponential boundary separating the relaxing (quantum) from the relaxed (classical) phase. The same relaxation behavior is observed for spinless fermions, which establishes the universality of such chaotic dynamics in interacting quantum systems. We also studied a quench protocol to weaker attractions and observed signals of quasi-equilibrium for truly long-ranged interactions.
@@ Acknowledgements
Appendix A Initial states
@@
Appendix B Stronger quench dynamics

### v2

@@ 7 Conclusions and Outlook
The present work addresses the nonequilibrium quench dynamics in a long-range interacting system. We investigated a generalized long-range interacting Bose gas where the strong repulsive interactions are suddenly quenched to strong attractive values, following the seminal observation of the exotic sTG gas in Haller’s experiment [66]. We presented a numerical study of the many-body dynamics for a wide range of interactions for both bosonic and fermionic systems. The usage of ab initio many-body methods facilitates a complete understanding of the exotic quenched phase in terms of fragmentation, delocalization and entropy production. We observed that the dynamics exhibited by the quench reveals new insights into the structure of quantum state relaxation.
Although the system eventually relaxes to a maximal entropy state, it passes through very complex phases of violent fragmentation and chaotic delocalization. The entropy production is determined by the number of configurations in the Hilbert space, which are saturated for all the numbers of orbitals that can be employed with current computational capabilities. The observation of unbounded entropy and chaotic spreading of the quenched gas in the available Hilbert space strongly suggests a behavior akin to that of a classical gas. We presented different measures of entropy for our discussion. However, we chose the occupation entropy as an order parameter to produce a parameter-time diagram that illustrates the relaxation process as a function of interaction length, covering the entire range of $4.00.5<\alpha<$ 4.0. We found an exponential boundary separating the relaxing (quantum) from the relaxed (classical) phase. The same relaxation behavior is observed for spinless fermions, which establishes the universality of such chaotic dynamics in interacting quantum systems. We also studied a quench protocol to weaker attractions and observed signals of dynamics slowdown and potential precursors to prethermalization for truly long-ranged interactions.
@@ Acknowledgements
Appendix A Appendix - Initial states
@@
Appendix B Appendix - Stronger quench dynamics
@@
Appendix C Appendix - Units
In this appendix, we discuss the units for the simulations presented in the main text. The system consists of $N=5$ bosons or fermions in an optical harmonic trap of frequency $\omega$, leading to the external potential $\hat{V}(x)=\frac{1}{2}\omega x^{2}$.
The units of our simulations are chosen as follows. We choose to set the unit of length as the inverse of the harmonic trapping frequency, i.e. $\bar{L}\equiv\sqrt{\hbar/(m\omega)}$. We then run simulations with 1024 gridpoints in an interval $x\in[-16\bar{L},16\bar{L}]$, giving a resolution of 0.03125 $\bar{L}$.
The unit of energy $\bar{E}$ in MCTDH-X is defined in terms of the unit of length as $\bar{E}\equiv\frac{\hbar^{2}}{m\bar{L}^{2}}$. By inserting our choice for the unit of length, we immediately see that the unit of energy corresponds to the quantized energy of the harmonic trap, i.e. $\bar{E}=\hbar\omega$. During the quench procedure, we change the strength of the long-range interactions from an initial value of $g_{d}=+100\bar{E}$ to a negative (attractive) value of between $g_{d}=-10\bar{E}$ and $g_{d}=-100\bar{E}$, as explained in the main text. Note that while strong, the initial repulsion is not enough for the particles to escape the harmonic confinement, as can be seen from Fig. 12.
The unit of time in MCTDH-X is also defined from the unit of length, as $\bar{t}\equiv\frac{m\hat{L}^{2}}{\hbar}$. Again inserting our definition of unit of length, we can simplify $\bar{t}=\frac{1}{\omega}$, i.e. the unit of time is the inverse frequency of the trap. In our simulations, we ran most time evolutions up until around $t\approx 1.2\bar{t}$ as this time scale is enough to probe the relaxation dynamics of the quench. Note that this is a relatively short time – just slightly above the time it takes for one complete oscillation in a harmonic potential. The fast dynamics we are generating is a consequence of the strong interactions employed.